You are here: Foswiki>Commissioning Web>Logs>26Jan013_Log (29 Jan 2013, JuanGuerra)Edit Attach

UT 20130126 AO#1

Observer:Esposito,S(Arcetri),,Puglisi,A(Arcetri),Pinna,E(Arcetri),JCGuerra(LBTO),JChristou(LBTO),
Telescope Operator: Steve..(LBTO)
SW Support: Cox,D (Tucson)
AO Support: JCGuerra(LBTO),JChristou(LBTO)
Telescope Support: Kevin N(LBTO)
Instrument Support: Tim S (LBTO)

Summary

AGW#2 is back to the telescope. RR is installed. The AGW#1 is in the lab working on the stage movements. Acquire source and close the loop using the AO procedures for day time.

Detail

The AGW#2 iws going to the telescope and the retro-reflector will be put on. The operation to install the AGW is about 4 hours plus the time to install the retro-reflector that it will take another hour.

About 1pm the systems is almost ready and to be tested.

At 3pm the system is ready to be tested and during the first test we have seen that the fibres connections were wrong, the WFS couldn't talk with the ASM, the workaround it was tro swaop the fibres in the tree house.

The Wavefront sensor is pwoer on and all its functionallity are running. Julian adn Juan Carlos were through the FLAO user procedures for day time work. Julian could get the spots in the IRTC and to close the loop.

After midnight other test were running using the system, closing the loop with different magnitudes and to get an first estimation of the performances. The BCU 47 didn't show any p[roblem working at low temperature.

The Arcetri team was working on with the AGW#1 in the lab, improving the stages functionallity.

-- JuanGuerra - 26 Jan 2013
Topic revision: r3 - 29 Jan 2013, JuanGuerra
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback